OpenStudy (anonymous):

if f(x)=x^2+4 and g(x)=2x+3, find f(g(-2))

OpenStudy (cwrw238):

g(-2) is found by plugging in x = -2 in g(x) then plug in g(-2) into f(x)

OpenStudy (cwrw238):

i'll do the first part for you g(-2) = 2(-2) + 3 = -4+3 = -1